# Energetic particles accelerated via turbulent magnetic reconnection in protoplanetary discs – I: ionisation rates

### v1

@@ 5.4 Limitations
First, on the acceleration model. We adopt a Fermi-like gain rate $\dot{E}=\alpha_{\rm acc}E$ and connect $\alpha_{\rm acc}$, $\tau_{\rm esc}$, and $\tau_{\rm inj}$ to a turbulent reconnection geometry. While these estimates are consistent with reconnection-in-turbulence theory and numerical trends, several efficiencies remain poorly constrained and depend on disc height and radius: (i) the partition of released magnetic energy into particles versus heat; (ii) the electron–proton sharing, future work will include injection and acceleration efficiencies $\eta_{e},\eta_{p}$; (iii) the fraction of turbulent power that feeds reconnection, set by the statistics and filling factor of current sheets; and (iv) the role of guide-field geometry and plasma $\beta_{p}$ in regulating $\alpha_{\rm acc},\ \tau_{\rm esc}$ and pitch-angle scattering. To make these dependencies explicit and to avoid over-constraining uncertain microphysics, we normalise the non-thermal component by the local energy budget through $\kappa$. The phenomenological parameter $\kappa$ thus absorbs uncertainties in the reconnection, particle/heat partition, and species sharing, allowing us to assess how the ionisation maps respond to realistic variations in the acceleration efficiency while remaining grounded in the turbulent–reconnection geometry. A physically grounded estimation of $\kappa$ will be provided in a future publication.
@@
Second, on particle transport. We model EP propagation using the CSDA, neglecting pitch–angle scattering. CSDA is reliable up to column densities of order $\sim 10^{25}\mathrm{cm^{-2}}$; beyond this depth (i.e. for protons above the pion–production threshold) a transport treatment including spatial diffusion and discrete losses is required (Padovani et al., 2018). In this setup we also approximate the magnetic field as locally vertical, omitting its toroidal and poloidal components. A magnetic field configuration, combining toroidal shear and poloidal lines, would induce helical guiding-centre motion, thereby modifying path lengths and redistributing EP flux laterally. Magnetic–field wandering would also lengthen particle path–lengths and generally reduce penetration depths (tending to shift EP–dominated regions upward in $Z/R$) and may inject particles at higher radii. In addition, field–line channelling could focus fluxes along preferred directions; capturing these effects would demand coupling transport to self–consistent MHD fields. Furthermore, in turbulent flows, non-thermal particles can further gain energy by stochastic acceleration in randomly moving MRI-generated magnetic perturbations (Sun and Bai, 2021). In this situation, non-thermal particles can also get (re)accelerated by Fermi second order process and shear acceleration. The particle maximum momentum $p_{\rm max}$ scales as the ratio of the local sound speed to light speed, $c_{s}/c$. Because of the small ratio $c_{s}/c$ it is clear that the acceleration effect is rather modest. Quantifying such re-acceleration is deferred to future work.
@@
Third, on time dependence and intermittency. We compute the non-thermal distribution at the end of a current-sheet lifetime and then average over acceleration sites on a stationary disc background. The accretion rate used is likewise steady and conservative ($\dot{M}=10^{-8}M_{\odot}\mathrm{yr}^{-1}$), i.e. it does not include episodic enhancements. In reality, accretion, reconnection, and turbulence are intermittent, and young discs show non-stationary accretion with frequent bursts. During such bursts, a larger fraction of the (temporarily increased) accretion power can be channelled into EPs, raising the ionisation and heating rates ($\zeta_{\rm EP}$, $\Gamma_{\rm EP}$), pushing EP influence deeper in the disc and driving time-variable chemical responses. Incorporating burst statistics and fully time-dependent transport and thermochemistry is a natural next step, beyond the steady-state framework adopted here.
@@
As a perspective three next steps are appearing. (1) Thermo-chemical feedback and observables: post-process with ProDiMo including $\zeta_{\rm EP}(R,Z)$ and $\Gamma_{\rm EP}(R,Z)$ to predict changes in abundances, temperatures, and synthetic observational ALMA/JWST diagnostics, and to identify discriminants between EP-, X-ray-, and CR-dominated regimes. (2) Calibration of efficiencies: constrain $\alpha_{\rm acc}$, $\tau_{\rm esc}$, $\tau_{\rm inj}$ and particle/heat partition from stratified-box simulations that measure reconnection statistics, and confront these with observational diagnostics. (3) Transport: couple EP propagation to 3D MHD fields to include field-line diffusion, and stochastic re-acceleration, and extend beyond the CSDA limit.
@@ 6 Conclusion
We investigated the ionisation produced by energetic particles (EPs) accelerated in situ by turbulence-induced magnetic reconnection in the atmospheres of protoplanetary discs. Building on a Fermi-like prescription for the energy gain and linking the acceleration/escape/injection timescales to a turbulent–reconnection geometry (Xu and Lazarian, 2023), we derived a non-thermal particle distribution with a steady-state spectral index $p=2.5$, an energy range bounded by a turbulence– and geometry–set cut-off, and a normalisation constrained by the fraction of accretion energy channelled to EPs through the parameter $\kappa$. Using a ProDiMo disc background, we computed EP propagation and the resulting ionisation and heating rate spatial distribution.
@@
    A new leading ionisation source. For $\kappa\gtrsim 10^{-2}$, EPs dominate the ionisation budget over stellar X-rays and Galactic cosmic rays (GCRs) in the inner disc (sub-AU to a few AU) at normalised height $Z/R\approx 0.1$ and in the disc atmosphere out to radii of a few tens of AU (see Figs. 8, 7). Even for $\kappa\sim 10^{-3}$, EPs contribute at the few-percent level locally—enough to affect chemistry and magnetic coupling. Below $10^{-3}$, they are minor.
@@
    Observability and locality as a test. Because EPs are generated locally they imprint spatially confined ionisation patterns distinct from globally distributed GCRs or stellar X-rays. This locality will provide testable signatures in emitting position of ion tracers (HCO+, N2H+) and mid-IR species (HCN, C2H2) in the same layers that JWST (especially MIRI) probes. Comparing these diagnostics against our maps could offer a way to test our model.
@@
In summary, turbulence-induced magnetic reconnection in disc atmospheres provides a physically motivated, in situ source of energetic particles that can reshape the ionisation structure of protoplanetary discs. Even when energetically subdominant, these particles penetrate and affect layers where other ionisers fade. When reconnection is efficient, they become a leading ionisation channel. Accounting for this component is therefore necessary for more realistic models of disc chemistry, dynamics, and their observational signatures. In turn, including this EP component in disc models will matter in the broader astrophysical context. In astrobiology, ionisation helps building (and destructing) complex organic molecules, sets where prebiotic species can form, and affects the interpretation of possible biosignatures in young planetary systems. More broadly, the same turbulence–reconnection framework can be adapted to accretion onto compact objects, where it may likewise contribute to the accretion–ejection dynamics.

### v2

@@ 5.4 Limitations
Our framework entails several simplifying assumptions. First, on the acceleration model. We adopted a Fermi-like gain rate, $\dot{E}=\alpha_{\rm acc}E$, and connected $\alpha_{\rm acc}$, $\tau_{\rm esc}$, and $\tau_{\rm inj}$ to a turbulent reconnection geometry. While these estimates are consistent with reconnection-in-turbulence theory and numerical trends, several efficiencies remain poorly constrained and depend on disc height and radius: (i) the partition of released magnetic energy into particles versus heat; (ii) the electron–proton sharing, for which future work will include injection and acceleration efficiencies, $\eta_{e},\eta_{p}$; (iii) the fraction of turbulent power that feeds reconnection, set by the statistics and filling factor of current sheets; and (iv) the role of guide-field geometry and plasma, $\beta_{p}$, in regulating $\alpha_{\rm acc},\ \tau_{\rm esc}$ and pitch-angle scattering. To make these dependencies explicit and to avoid over-constraining uncertain microphysics, we normalised the non-thermal component by the local energy budget through $\kappa$. The phenomenological parameter $\kappa$ thus absorbs uncertainties in the reconnection, particle and heat partition, and species sharing, allowing us to assess how the ionisation maps respond to realistic variations in the acceleration efficiency, while remaining grounded in the turbulent–reconnection geometry. A physically grounded estimation of $\kappa$ will be provided in a future publication.
@@
Second, on particle transport, we modelled EP propagation using the CSDA, neglecting pitch–angle scattering. CSDA is reliable up to column densities of order $\sim 10^{25}\mathrm{cm^{-2}}$; beyond this depth (i.e. for protons above the pion–production threshold) a transport treatment including spatial diffusion and discrete losses is required (Padovani et al., 2018). In this set-up we also approximate the magnetic field as locally vertical, omitting its toroidal and poloidal components. A magnetic field configuration, combining toroidal shear and poloidal lines, would induce helical guiding-centre motion, thereby modifying path lengths and redistributing EP flux laterally. Magnetic field wandering would also lengthen particle path–lengths and generally reduce penetration depths (tending to shift EP–dominated regions upwards in $Z/R$), and may inject particles at higher radii. In addition, field–line channelling could focus fluxes in preferred directions; capturing these effects would demand coupling transport to self-consistent MHD fields. Furthermore, in turbulent flows, non-thermal particles can further gain energy by stochastic acceleration in randomly moving MRI-generated magnetic perturbations (Sun and Bai, 2021). In this situation, non-thermal particles can also get (re)accelerated by Fermi second order process and shear acceleration. The particle maximum momentum, $p_{\rm max}$, scales as the ratio of the local sound speed to light speed, $c_{s}/c$. Because of the small ratio $c_{s}/c$, it is clear that the acceleration effect is rather modest. Quantifying such re-acceleration is deferred to future work.
@@
Third, on time dependence and intermittency, we computed the non-thermal distribution at the end of a current-sheet lifetime and then averaged over acceleration sites on a stationary disc background. The accretion rate used is likewise steady and conservative ($\dot{M}=10^{-8}M_{\odot}\mathrm{yr}^{-1}$), i.e. it does not include episodic enhancements. In reality, accretion, reconnection, and turbulence are intermittent, and young discs show non-stationary accretion with frequent bursts. During such bursts, a larger fraction of the (temporarily increased) accretion power can be channelled into EPs, raising the ionisation and heating rates ($\zeta_{\rm EP}$, $\Gamma_{\rm EP}$), pushing EP influence deeper in the disc and driving time-variable chemical responses. Incorporating burst statistics and fully time-dependent transport and thermochemistry is a natural next step, beyond the steady-state framework adopted here.
@@
As a perspective three next steps are appearing. (1) Thermo-chemical feedback and observables: post-process with ProDiMo including $\zeta_{\rm EP}(R,Z)$ and $\Gamma_{\rm EP}(R,Z)$ to predict changes in abundances, temperatures, and synthetic observational ALMA/JWST diagnostics, and to identify discriminants between EP-, X-ray-, and CR-dominated regimes. (2) Calibration of efficiencies: constrain $\alpha_{\rm acc}$, $\tau_{\rm esc}$, $\tau_{\rm inj}$ and particle and heat partition from stratified-box simulations that measure reconnection statistics, and confront these with observational diagnostics. (3) Transport: couple EP propagation to 3D MHD fields to include field-line diffusion, and stochastic re-acceleration, and extend beyond the CSDA limit.
@@ 6 Conclusion
We investigated the ionisation produced by EPs accelerated in situ by turbulence-induced magnetic reconnection in the atmospheres of protoplanetary discs. Building on a Fermi-like prescription for the energy gain and linking the acceleration, escape and injection timescales to a turbulent–reconnection geometry (Xu and Lazarian, 2023), we derived a non-thermal particle distribution with a steady-state spectral index $p=2.5$, an energy range bounded by a turbulence– and geometry–set cut-off, and a normalisation constrained by the fraction of accretion energy channelled to EPs through the parameter $\kappa$. Using a ProDiMo disc background, we computed EP propagation and the resulting ionisation and heating rate spatial distribution.
@@
    A new leading ionisation source. For $\kappa\gtrsim 10^{-2}$, EPs dominate the ionisation budget over stellar X-rays and GCRs in the inner disc (sub-astronomical units to a few astronomical units) at normalised height $Z/R\approx 0.1$ and in the disc atmosphere out to radii of a few tens of astronomical units (see Figs. 8, 7). Even for $\kappa\sim 10^{-3}$, EPs contribute at the level of a few percent locally – enough to affect chemistry and magnetic coupling. Below $10^{-3}$, they are minor.
@@
    Observability and locality as a test. Because EPs are generated locally, they imprint spatially confined ionisation patterns distinct from globally distributed GCRs or stellar X-rays. This locality will provide testable signatures in emitting position of ion tracers (HCO+, N2H+) and mid-IR species (HCN, C2H2) in the same layers that JWST (especially MIRI) probes. Comparing these diagnostics against our maps could offer a way to test our model.
@@
In summary, turbulence-induced magnetic reconnection in disc atmospheres provides a physically motivated, in situ source of EPs that can reshape the ionisation structure of protoplanetary discs. Even when energetically subdominant, these particles penetrate and affect layers where other ionisers fade. When reconnection is efficient, they become a leading ionisation channel. Accounting for this component is therefore necessary for more realistic models of disc chemistry, dynamics, and their observational signatures. In turn, including this EP component in disc models will matter in the broader astrophysical context. In astrobiology, ionisation helps building (and destructing) complex organic molecules, sets where prebiotic species can form, and affects the interpretation of possible biosignatures in young planetary systems. More broadly, the same turbulence–reconnection framework can be adapted to accretion onto compact objects, where it may likewise contribute to the accretion–ejection dynamics.